ON BIVARIATE LACK OF MEMORY PROPERTY AND A NEW DEFINITION

DILIP ROY

Department of Business Administration, Burdwan University, Burdwan–713104, WB, India

(Received February 22, 2000; revised October 12, 2000)

Abstract. The earlier concept of bivariate lack of memory property has been examined and an alternative concept of bivariate lack of memory property has been presented along with a set of equivalent conditions. In this process, the link relations known in the literature on univariate lack of memory property have been extended to the bivariate set-up. A multivariate extension of the lack of memory property has also been proposed.

Key words and phrases: Lack of memory property, failure rate, mean residual life, coefficient of variation of residual life, bivariate extension.